REU: REVISED SCHEDULE FOR WEEKS SEVEN AND EIGHT

BABAI’S AFTERNOON TALKS ARE IN R276 MAY’S MORNING TALKS ARE IN R276 OTHER TALKS ARE IN ECKHART 206

Ex–YSP participants are expected to arrange mentoring sessions with graduate students and problem–solving group sessions between 9:30 and 4:00. This is time free for you to work on problems.

Date  SESAME  Babai  Bader  Kerr  May
Mon, July 31  9–12, 1–4 3:00 – 4:45  5:00 – 6:00 11:00-12:00
Tue, August 1 9–12, 1–4  4:30 – 6:00
Wed, August 2 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 10:30-12:00
Thu, August 3 9–12, 1–4  4:30 – 6:00
Fri, August 4 9–12, 1–4 3:00 – 4:45 5:00 – 6:00 10:30-12:00
Mon, August 7 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 10:30-12:00
Tue, August 8 9–12, 1–4  4:30–6:00
Wed, August 9 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 10:30-12:00
Thu, August 10 9–12, 1–4  4:30–6:00
Fri, August 11 9–12, 1–4 3:00 – 4:45  5:00 – 6:00 10:30-12:00

Here is the schedule in summary form:

SEVENTH WEEK:

SESAME: 9:00 – 12:00 and 1:00 – 4:00 M-F Babai (Discrete): 3:00 – 4:45 MWF (SESAME participants: come on in) Bader (Geometry): 4:30 – 6:00 TTh Kerr (Algebraic geometry): 5:00 – 6:00 MWF May (For those who can come): 10:30 – 12:00 MWF

EIGHTH WEEK:

SESAME: 9:00 – 12:00 and 1:00 – 4:00 M-F Babai (Discrete): 3:00 – 4:45 MWF (SESAME participants: come on in) Bader (Geometry): 4:30 – 6:00 TTh Kerr (Algebraic geometry): 5:00 – 6:00 MWF May (For those who can come): 10:30 – 12:00 MWF

Expected, August 9-11:

WEDNESDAY, 6:30 pm: Student presentations and dinner THURSDAY, 6:30 pm: Student presentations and dinner FRIDAY, 6:30: Possibly more student presentations